# Relational Representation Distillation

### v3

@@ 4.1 Results on CIFAR-100
Table 1 and Table 2 present the top-1 accuracies of student networks trained using different distillation techniques across various teacher-student architectural pairings. Table 1 examines pairings where both student and teacher models share similar architectural styles, while Table 2 focuses on cross-architecture distillations. Our proposed loss consistently outperforms the conventional KD technique, as indicated by the green upward arrows ($\uparrow$). While the standalone performance of our method is comparable to CRD, its integration with KD not only achieves higher accuracies but in some cases, surpasses the performance of the teacher networks, such as in the distillation of WRN-40-2 to ShuffleNet-v1. The enhanced performance of our distillation method can be credited to multiple factors that collectively improve the transfer of knowledge from the teacher to the student model. Our approach uses a unique loss function that complements KD’s primary focus on matching the softened output logits of the teacher and student. We introduce an additional layer of representational alignment that ensures not only the final outputs but also the intermediate feature representations of the student closely match those of the teacher. This dual focus allows the student model to mimic the teacher’s outputs and develop more robust and generalizable internal representations.
@@
FitNet [46] 73.58 ($\downarrow$) 72.24 ($\downarrow$) 69.21 ($\downarrow$) 68.99 ($\downarrow$) 71.06 ($\downarrow$) 73.50 ($\downarrow$) 71.02 ($\downarrow$)
AT [59] 74.08 ($\downarrow$) 72.77 ($\downarrow$) 70.55 ($\downarrow$) 70.22 ($\downarrow$) 72.31 ($\downarrow$) 73.44 ($\downarrow$) 71.43 ($\downarrow$)
SP [52] 73.83 ($\downarrow$) 72.43 ($\downarrow$) 69.67 ($\downarrow$) 70.04 ($\downarrow$) 72.69 ($\downarrow$) 72.94 ($\downarrow$) 72.68 ($\downarrow$)
CC [44] 73.56 ($\downarrow$) 72.21 ($\downarrow$) 69.63 ($\downarrow$) 69.48 ($\downarrow$) 71.48 ($\downarrow$) 72.97 ($\downarrow$) 70.81 ($\downarrow$)
VID [1] 74.11 ($\downarrow$) 73.30 ($\downarrow$) 70.38 ($\downarrow$) 70.16 ($\downarrow$) 72.61 ($\downarrow$) 73.09 ($\downarrow$) 71.23 ($\downarrow$)
RKD [41] 73.35 ($\downarrow$) 72.22 ($\downarrow$) 69.61 ($\downarrow$) 69.25 ($\downarrow$) 71.82 ($\downarrow$) 71.90 ($\downarrow$) 71.48 ($\downarrow$)
PKT [42] 74.54 ($\downarrow$) 73.45 ($\downarrow$) 70.34 ($\downarrow$) 70.25 ($\downarrow$) 72.61 ($\downarrow$) 73.64 ($\uparrow$) 72.88 ($\downarrow$)
AB [21] 72.50 ($\downarrow$) 72.38 ($\downarrow$) 69.47 ($\downarrow$) 69.53 ($\downarrow$) 70.98 ($\downarrow$) 73.17 ($\downarrow$) 70.94 ($\downarrow$)
FT [27] 73.25 ($\downarrow$) 71.59 ($\downarrow$) 69.84 ($\downarrow$) 70.22 ($\downarrow$) 72.37 ($\downarrow$) 72.86 ($\downarrow$) 70.58 ($\downarrow$)
FSP [58] 72.91 ($\downarrow$) n/a 69.95 ($\downarrow$) 70.11 ($\downarrow$) 71.89 ($\downarrow$) 72.62 ($\downarrow$) 70.33 ($\downarrow$)
NST [24] 73.68 ($\downarrow$) 72.24 ($\downarrow$) 69.60 ($\downarrow$) 69.53 ($\downarrow$) 71.96 ($\downarrow$) 73.30 ($\downarrow$) 71.53 ($\downarrow$)
CRD [51] 75.48 ($\uparrow$) 74.14 ($\uparrow$) 71.16 ($\uparrow$) 71.46 ($\uparrow$) 73.48 ($\uparrow$) 75.51 ($\uparrow$) 73.94 ($\uparrow$)
CRD+KD [51] 75.64 ($\uparrow$) 74.38 ($\uparrow$) 71.63 ($\uparrow$) 71.56 ($\uparrow$) 73.75 ($\uparrow$) 75.46 ($\uparrow$) 74.29 ($\uparrow$)
OFD [20] 75.24 ($\uparrow$) 74.33 ($\uparrow$) 70.38 ($\uparrow$) n/a 73.23 ($\uparrow$) 74.95 ($\uparrow$) 73.95 ($\uparrow$)
WSLD [64] n/a 73.74 ($\uparrow$) 71.53 ($\uparrow$) n/a 73.36 ($\uparrow$) 74.79 ($\uparrow$) n/a
IPWD [38] n/a 74.64 ($\uparrow$) 71.32 ($\uparrow$) n/a 73.91 ($\uparrow$) 76.03 ($\uparrow$) n/a
RRD (ours) 75.33 ($\uparrow$) 73.55 ($\uparrow$) 70.71 ($\uparrow$) 70.72 ($\uparrow$) 73.10 ($\uparrow$) 74.48 ($\uparrow$) 73.99 ($\uparrow$)
RRD+KD (ours) 75.66 ($\uparrow$) 74.67 ($\uparrow$) 72.19 ($\uparrow$) 71.74 ($\uparrow$) 73.54 ($\uparrow$) 75.08 ($\uparrow$) 74.32 ($\uparrow$)
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column.
@@
FitNet [46] 64.14 ($\downarrow$) 63.16 ($\downarrow$) 70.69 ($\downarrow$) 73.59 ($\downarrow$) 73.54 ($\downarrow$) 73.73 ($\downarrow$)
AT [59] 59.40 ($\downarrow$) 58.58 ($\downarrow$) 71.84 ($\downarrow$) 71.73 ($\downarrow$) 72.73 ($\downarrow$) 73.32 ($\downarrow$)
SP [52] 66.30 ($\downarrow$) 68.08 ($\downarrow$) 73.34 ($\downarrow$) 73.48 ($\downarrow$) 74.56 ($\uparrow$) 74.52 ($\downarrow$)
CC [44] 64.86 ($\downarrow$) 65.43 ($\downarrow$) 70.25 ($\downarrow$) 71.14 ($\downarrow$) 71.29 ($\downarrow$) 71.38 ($\downarrow$)
VID [1] 65.56 ($\downarrow$) 67.57 ($\downarrow$) 70.30 ($\downarrow$) 73.38 ($\downarrow$) 73.40 ($\downarrow$) 73.61 ($\downarrow$)
RKD [41] 64.52 ($\downarrow$) 64.43 ($\downarrow$) 71.50 ($\downarrow$) 72.28 ($\downarrow$) 73.21 ($\downarrow$) 72.21 ($\downarrow$)
PKT [42] 67.13 ($\downarrow$) 66.52 ($\downarrow$) 73.01 ($\downarrow$) 74.10 ($\uparrow$) 74.69 ($\uparrow$) 73.89 ($\downarrow$)
AB [21] 66.06 ($\downarrow$) 67.20 ($\downarrow$) 70.65 ($\downarrow$) 73.55 ($\downarrow$) 74.31 ($\downarrow$) 73.34 ($\downarrow$)
FT [27] 61.78 ($\downarrow$) 60.99 ($\downarrow$) 70.29 ($\downarrow$) 71.75 ($\downarrow$) 72.50 ($\downarrow$) 72.03 ($\downarrow$)
NST [24] 58.16 ($\downarrow$) 64.96 ($\downarrow$) 71.28 ($\downarrow$) 74.12 ($\uparrow$) 74.68 ($\uparrow$) 76.09 ($\uparrow$)
CRD [51] 69.73 ($\uparrow$) 69.11 ($\uparrow$) 74.30 ($\uparrow$) 75.11 ($\uparrow$) 75.65 ($\uparrow$) 76.05 ($\uparrow$)
CRD+KD [51] 69.94 ($\uparrow$) 69.54 ($\uparrow$) 74.58 ($\uparrow$) 75.12 ($\uparrow$) 76.05 ($\uparrow$) 76.27 ($\uparrow$)
OFD [20] 69.48 ($\uparrow$) 69.04 ($\uparrow$) n/a 75.98 ($\uparrow$) 76.82 ($\uparrow$) 75.85 ($\uparrow$)
WSLD [64] n/a 68.79 ($\uparrow$) 73.80 ($\uparrow$) 75.09 ($\uparrow$) n/a 75.23 ($\uparrow$)
RRD (ours) 67.93 ($\uparrow$) 68.84 ($\uparrow$) 74.01 ($\uparrow$) 74.11 ($\uparrow$) 74.80 ($\uparrow$) 74.98 ($\uparrow$)
RRD+KD (ours) 69.98 ($\uparrow$) 69.13 ($\uparrow$) 74.26 ($\uparrow$) 75.18 ($\uparrow$) 76.83 ($\uparrow$) 76.31 ($\uparrow$)
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column.
@@ 4.2 Results on ImageNet
AT [59] 71.03 ($\uparrow$) 71.18 ($\downarrow$) 70.18 ($\downarrow$)
SP [52] 70.62 ($\uparrow$) 71.08 ($\downarrow$) n/a
CC [44] 69.96 ($\downarrow$) n/a n/a
VID [1] n/a 71.11 ($\downarrow$) n/a
RKD [41] 70.40 ($\downarrow$) n/a 68.50 ($\downarrow$)
AB [21] n/a n/a 68.89 ($\downarrow$)
FT [27] n/a n/a 69.88 ($\downarrow$)
FSP [58] 70.58 ($\downarrow$) n/a n/a
NST [24] 70.29 ($\downarrow$) n/a n/a
CRD [51] 71.17 ($\uparrow$) 71.25 ($\downarrow$) 69.07 ($\downarrow$)
OFD [20] 71.03 ($\uparrow$) n/a 71.33 ($\uparrow$)
WSLD [64] 72.04 ($\uparrow$) n/a 71.52 ($\uparrow$)
IPWD [38] 71.88 ($\uparrow$) n/a 72.65 ($\uparrow$)
RRD (ours) 71.22 ($\uparrow$) 71.33 ($\uparrow$) 70.66 ($\uparrow$)
RRD+KD (ours) 71.40 ($\uparrow$) 71.51 ($\uparrow$) 71.83 ($\uparrow$)
Table 3: Test top-1 (%) on ImageNet validation test using various distillation methods. The table compares students and teachers of the same and different architecture. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column while underlined values mark the second best.
4.3 Capturing Inter-class Correlations
Cross-entropy loss overlooks the relationships among class logits in a teacher network, often resulting in less effective knowledge transfer. Distillation techniques that use "soft targets", such as those described by [22], have successfully captured these relationships, improving student model performance. Figure 2 assesses the effectiveness of different distillation methods on the CIFAR-100 KD task using WRN-40-2 as the teacher and WRN-40-1 as the student. We compare students trained without distillation, with attention transfer [59], with KL divergence [22], and with our proposed RRD method. Our findings show that RRD achieves close alignment between teacher and student logits, as evidenced by reduced differences in their correlation matrices. While RRD does not match CRD [51] in terms of exact correlation alignment, it significantly enhances learning efficiency and reduces error rates. The smaller discrepancies between teacher and student logits indicate that the RRD objective captures a substantial portion of the correlation structure in the logits, resulting in lower error rates, though CRD achieves a slightly closer match.
(a) Student: vanilla
(b) Student: AT
(c) Student: KD
(d) Student: RRD (ours)
Figure 2: Comparison of correlation matrix differences between teacher and student logits across various distillation methods on the CIFAR-100 task. Subfigures show results for (a) students trained without distillation, (b) with attention transfer [59], (c) with KL divergence [22], and (d) with our RRD method, highlighting better matching between student’s and teacher’s correlations. Results have been re-implemented according to [51].
@@ 4.4 Transferability of Representations
Our study investigates how knowledge is transferred from a larger teacher network (WRN-40-2) to a smaller student network (WRN-16-2), aiming to develop versatile representations suitable for a range of tasks and datasets. We apply this technique by having the student network either learn directly from the CIFAR-100 dataset or via distillation. In this setup, the student network is employed as a static feature extractor for images from the STL-10 and TIN-200 datasets, both adjusted to a $32\times 32$ resolution. We evaluate the adaptability of these features by training a linear classifier on the final feature layer to conduct classifications with 10 categories for STL-10 and 200 categories for TIN-200. We document the impact of various distillation approaches on the transferability of these features in Table 4. Our findings reveal that, except for the FitNet method, all distillation techniques significantly improve the feature transferability on both datasets. Notably, while the teacher network achieves the highest performance on CIFAR-100, its features show the least effective transfer, likely due to training data specificity. Conversely, the student network using a combination of RRD and KD distillation techniques not only equates to the teacher’s CIFAR-100 performance but also surpasses it in transferability, showing improvements of 3.1% and 3.1% in STL-10 and TIN-200 respectively.
@@
Table 4: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and TIN-200 datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200). Results adapted from [51]. The values in bold indicate the maximum of each row.
@@ 4.5 Visualization of t-SNE Embeddings
Figure 3: t-SNE visualizations of embeddings from the teacher network and student networks trained using different distillation techniques on the first 10 classes of the CIFAR-100 dataset.
@@ 4.6 Ablation Study
Number of negatives $\bm{M}$. We validated different values for $M$: 256, 1024, 2048, 4096, 8192, 16384, and 32768. As shown in Figure 4(a), increasing $M$ leads to improved performance. However, the difference in error rate between $M=4096$ and $M=16384$ is less 0.5%. Therefore, we use $M=16384$ for reporting the accuracy, while in practice lower $M$ should suffice. Going beyond $M=16384$ proves to harms performance.
Temperature $\bm{\tau_{s}}$ and $\bm{\tau_{t}}$. We varied $\tau_{s}$ and $\tau_{t}$, considering all the permutations of the numbers 0.04, 0.07, and 0.2 taken two at a time, considering that order matters. As Figure 4(b) illustrates, both extremely high or low temperatures lead to a sub-optimal solution. Also, as expected, a lower $\tau_{s}$ than $\tau_{t}$ leads to better performance. This improvement is attributed to the sharper predictions from the student model with a lower temperature, resulting in more confident and distinct class probabilities. This sharpness helps the student model align better with the teacher’s guidance, enhance generalization, provide a more informative learning signal, and reduce overfitting. The sweep spot lies at $\tau_{s}=0.04$ and $\tau_{t}=0.07$ (red area of Figure 4(b)).
Loss coefficient $\bm{\beta}$. We varied $\beta$ from 0.1 to 100. As Figure 4(c) illustrates, both extremely high or low $\beta$ lead to a sub-optimal solution. In general, $\beta$ between 1 and 2 works well on CIFAR-100.
Computational cost. The RRD loss method introduces an additional 5.24 MFLOPs to the computational workload of a ResNet-50 model. This represents approximately 0.262% of the original 2 GFLOPs computational baseline. Despite the added computational complexity, the impact on training time is minimal, allowing us to maintain efficiency in practical training scenarios such as on CIFAR-100. The memory bank for RRD stores all 128-dimensional features of up to 16,384 CIFAR-100 images, consuming about 8 MB of GPU memory.

### v4

@@ 4.2 Results on CIFAR-100
Table 1 and Table 2 present the top-1 accuracies of student networks trained using different distillation techniques across various teacher-student architectural pairings. Table 1 examines pairings where both student and teacher models share similar architectural styles, while Table 2 focuses on cross-architecture distillations. Our proposed loss consistently outperforms the conventional KD technique. While the standalone performance of our method is comparable to CRD, its integration with KD not only achieves higher accuracies but in some cases, surpasses the performance of the teacher networks, such as in the distillation of WRN-40-2 to ShuffleNet-v1. The enhanced performance of our distillation method can be credited to multiple factors that collectively improve the transfer of knowledge from the teacher to the student model. Our approach uses a unique loss function that complements KD’s primary focus on matching the softened output logits of the teacher and student. We introduce an additional layer of representational alignment that ensures not only the final outputs but also the intermediate feature representations of the student closely match those of the teacher. This dual focus allows the student model to mimic the teacher’s outputs and develop more robust and generalizable internal representations.
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. The values in bold indicate the maximum of each column and underlined values mark the second best.
@@
FitNet [42] 73.58 72.24 69.21 68.99 71.06 73.50 71.02
AT [54] 74.08 72.77 70.55 70.22 72.31 73.44 71.43
SP [47] 73.83 72.43 69.67 70.04 72.69 72.94 72.68
CC [40] 73.56 72.21 69.63 69.48 71.48 72.97 70.81
VID [1] 74.11 73.30 70.38 70.16 72.61 73.09 71.23
RKD [37] 73.35 72.22 69.61 69.25 71.82 71.90 71.48
PKT [38] 74.54 73.45 70.34 70.25 72.61 73.64 72.88
AB [19] 72.50 72.38 69.47 69.53 70.98 73.17 70.94
FT [25] 73.25 71.59 69.84 70.22 72.37 72.86 70.58
FSP [53] 72.91 n/a 69.95 70.11 71.89 72.62 70.33
NST [22] 73.68 72.24 69.60 69.53 71.96 73.30 71.53
CRD [46] 75.48 74.14 71.16 71.46 73.48 75.51 73.94
CRD+KD [46] 75.64 74.38 71.63 71.56 73.75 75.46 74.29
OFD [18] 75.24 74.33 70.38 n/a 73.23 74.95 73.95
WSLD [59] n/a 73.74 71.53 n/a 73.36 74.79 n/a
IPWD [34] n/a 74.64 71.32 n/a 73.91 76.03 n/a
RRD (ours) 75.33 73.55 70.71 70.72 73.10 74.48 73.99
RRD+KD (ours) 75.66 74.67 72.19 71.74 73.54 75.08 74.32
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. The values in bold indicate the maximum of each column and underlined values mark the second best.
@@
FitNet [42] 64.14 63.16 70.69 73.59 73.54 73.73
AT [54] 59.40 58.58 71.84 71.73 72.73 73.32
SP [47] 66.30 68.08 73.34 73.48 74.56 74.52
CC [40] 64.86 65.43 70.25 71.14 71.29 71.38
VID [1] 65.56 67.57 70.30 73.38 73.40 73.61
RKD [37] 64.52 64.43 71.50 72.28 73.21 72.21
PKT [38] 67.13 66.52 73.01 74.10 74.69 73.89
AB [19] 66.06 67.20 70.65 73.55 74.31 73.34
FT [25] 61.78 60.99 70.29 71.75 72.50 72.03
NST [22] 58.16 64.96 71.28 74.12 74.68 76.09
CRD [46] 69.73 69.11 74.30 75.11 75.65 76.05
CRD+KD [46] 69.94 69.54 74.97 75.12 76.05 76.27
OFD [18] 69.48 69.04 n/a 75.98 76.82 75.85
WSLD [59] n/a 68.79 73.80 75.09 n/a 75.23
IPWD [34] n/a 70.25 74.58 76.03 n/a 76.44
RRD (ours) 67.93 68.84 74.01 74.11 74.80 74.98
RRD+KD (ours) 69.98 69.13 74.26 76.83 76.83 76.44
@@ 4.3 Results on ImageNet
Table 3: Test top-1 (%) on ImageNet validation set using various distillation methods. The table compares students and teachers of the same and different architecture. The values in bold indicate the maximum of each column while underlined values mark the second best.
@@
AT [54] 71.03 71.18 70.18
SP [47] 70.62 71.08 n/a
CC [40] 69.96 n/a n/a
VID [1] n/a 71.11 n/a
RKD [37] 70.40 n/a 68.50
AB [19] n/a n/a 68.89
FT [25] n/a n/a 69.88
FSP [53] 70.58 n/a n/a
NST [22] 70.29 n/a n/a
CRD [46] 71.17 71.25 69.07
OFD [18] 71.03 n/a 71.33
WSLD [59] 72.04 n/a 71.52
IPWD [34] 71.88 n/a 72.65
RRD (ours) 71.22 71.33 70.66
RRD+KD (ours) 71.40 71.51 71.83
@@ 4.4 Transferability of Representations
Our study investigates knowledge transfer from a larger teacher network (WRN-40-2) to a smaller student network (WRN-16-2), aiming to develop versatile representations for various tasks and datasets. We apply this technique by having the student network learn directly from CIFAR-100 or via distillation. The student network is employed as a static feature extractor for STL-10 and Tiny ImageNet images, adjusted to $32\times 32$ resolution. We evaluate feature adaptability by training a linear classifier on the final feature layer for classifications with 10 categories (STL-10) and 200 categories (Tiny ImageNet). We document the impact of distillation approaches on feature transferability in Table 4. Our findings reveal that, except for FitNet, all distillation techniques significantly improve feature transferability on both datasets. Notably, while the teacher network achieves highest performance on CIFAR-100, its features show least effective transfer, likely due to training data specificity. Conversely, the student network using RRD and KD distillation not only equates to the teacher’s CIFAR-100 performance but also surpasses it in transferability, showing improvements of 3.1% in STL-10 and Tiny ImageNet.
Table 4: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and Tiny ImageNet datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (Tiny ImageNet). Results adapted from [46]. The values in bold indicate the maximum of each row.
@@
4.5 Capturing Inter-class Correlations
Cross-entropy loss overlooks the relationships among class logits in a teacher network, often resulting in less effective knowledge transfer. Distillation techniques that use ”soft targets”, such as those described by [20], have successfully captured these relationships, improving student model performance. Figure 2 assesses the effectiveness of different distillation methods on the CIFAR-100 KD task using WRN-40-2 as the teacher and WRN-40-1 as the student. We compare students trained without distillation, with attention transfer [54], with KL divergence [20], and with our proposed RRD method. Our findings show that RRD achieves close alignment between teacher and student logits, as evidenced by reduced differences in their correlation matrices. While RRD does not match CRD [46] in terms of exact correlation alignment, it significantly enhances learning efficiency and reduces error rates. The smaller discrepancies between teacher and student logits indicate that the RRD objective captures a substantial portion of the correlation structure in the logits, resulting in lower error rates, though CRD achieves a slightly closer match.
(a) Student: vanilla
(b) Student: KD [20]
(c) Student: AT [54]
(d) Student: RRD (ours)
Figure 2: Comparison of correlation matrix differences between teacher and student logits across various distillation methods on the CIFAR-100 task. Subfigures show results for (a) students trained without distillation, (b) with KL divergence [20], (c) with attention transfer (AT) [54], and (d) with our RRD method, highlighting better matching between student’s and teacher’s correlations. Results have been re-implemented according to [46].
@@ 4.6 Visualization of t-SNE Embeddings
Figure 3: t-SNE visualizations of embeddings from the teacher network and student networks trained using different distillation techniques on the first 10 classes of the CIFAR-100 dataset: (a) teacher, (b) student trained without distillation, (c) student trained with attention transfer (AT) [54], and (d) student trained with RRD.
@@ 4.7 Ablation Study
Ablation on number of negatives $\bm{M}$.
We validated different values for $M$: 256, 1024, 2048, 4096, 8192, 16384, and 32768. As shown in Figure 4(a), increasing $M$ leads to improved performance. However, the difference in error rate between $M=4096$ and $M=16384$ is less 0.5%. Therefore, we use $M=16384$ for reporting the accuracy, while in practice lower $M$ should suffice. Going beyond $M=16384$ proves to harms performance.
Ablation on temperatures $\bm{\tau_{s}}$ and $\bm{\tau_{t}}$.
We varied $\tau_{s}$ and $\tau_{t}$, considering all the permutations of the numbers 0.04, 0.07, and 0.2 taken two at a time, considering that order matters. As Figure 4(b) illustrates, both extremely high or low temperatures lead to a sub-optimal solution. Also, as expected, a lower $\tau_{s}$ than $\tau_{t}$ leads to better performance. This improvement is attributed to the sharper predictions from the student model with a lower temperature, resulting in more confident and distinct class probabilities. This sharpness helps the student model align better with the teacher’s guidance, enhance generalization, provide a more informative learning signal, and reduce overfitting. The sweep spot lies at $\tau_{s}=0.04$ and $\tau_{t}=0.07$ (red area of Figure 4(b)).
Ablation on loss coefficient $\bm{\beta}$.
We varied $\beta$ from 0.1 to 100. As Figure 4(c) illustrates, both extremely high or low $\beta$ lead to a sub-optimal solution. In general, $\beta$ between 1 and 2 works well on CIFAR-100, as it balances the KD loss effectively without overwhelming other components of the training objective.